mikewwe13:

The function f(x)=80(1.5)^x models a bacteria population after x hours. How does the average rate of change between Hour 4 and Hour 8 compare to the average rate of change between Hour 0 and Hour 4? A. The average rate of change is 4 times as fast. B. The average rate of change is 2.25 times as fast. C. The average rate of change is 1.5 times as fast. D. The average rate of change is 5.0625 times as fast.

mikewwe13:

@Vocaloid

sillybilly123:

Because \(\Delta x_{1,2} = 4\) we only need to compare \(\Delta f_1\) vs \(\Delta f_2\) And: \( \dfrac{\Delta f_2}{\Delta f_1} = \dfrac{ 1.5^8 - 1.5^4}{1.5^4 - 1.5^0} = (D)\) if you follow that through....

Vocaloid:

@mikewwe13 if you're having trouble understanding some of that notation let me know ^^ they might not have learned about deltas yet